OpenStudy (anonymous):

Radio Stations The approximate number of radio stations on the air for selected years from 1950 to 2000 is shown in the table. x (year) 1950 1960 1970 y (stations) 2800 4100 6800 x (year) 1980 1990 2000 y (stations) 8600 10,800 12,600 Estimation Discuss ways to estimate the number of radio stations on the air in 1975. Compare yourestimates with the actual value of 7700 stations. Repeat this estimate for 1985 and compare it to the actual value of 10,400. Discuss your results.

OpenStudy (anonymous):

Is it mathematics or space question ? !!!!! I like your teacher .. sorry i wish 2 help . good luck

OpenStudy (dumbcow):

you could do a linear regression: y = 202x + 2552, where x is years after 1950

OpenStudy (anonymous):

dumbcow is right. You can try to fit a line to the data, and then use that line to predict stations for 1975 and 1985

OpenStudy (anonymous):

Please where from that equation y = 202x + 2552

OpenStudy (ash2326):

for 1975,x= 1975-1950 so x =25 y=202x+25 y=5050+25 y=5075, so there is a large gap from actual value of 7700 for 1985 x=1985-1950=35 y=202x+25 y=202*35+25=7105 so there is a large variation from actual value of 10, 400

OpenStudy (dumbcow):

i used a graphing calculator...Excel could also do it for you

OpenStudy (anonymous):

Please,Dumbcow for 1975 and 1950,that means I will use 25 and for 1985 and 1950 I will use 35 for the values of x?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

Please Dumbcow can you tell me how u got the equation,more especially with the 2552

OpenStudy (dumbcow):

its in the excel sheet i attached...make a chart, add a trend line and Excel does the regression for you. Thats how i got the equation

OpenStudy (dumbcow):

For graphing calculators, insert data points into 2 lists then choose LinReg from "STAT"->"CALC"